Angela M. Reese

2763 South Norfolk Street #111, San Mateo, CA 94403 amreese85@gmail.com (513) 341-5840

Education

• Stephen T. Badin High School — Hamilton, OH

Student: 1999 – 2003 – Graduated with over 80 hours of volunteer community service.

Experience

• Amy’s Hallmark — San Mateo, CA

Sales Associate: January 2012 – Present – Supervisor: Laurel Overstreet

– Assist customers in merchandise selection and purchase.

– Organize and maintain product displays and placement.

• Humana RightSource Rx — West Chester, OH

Certified Pharmacy Technician: November 2009 – October 2011 – Supervisor: Chantel Jones

– Resolve issues with prescriptions by contacting doctors or their staff.

– Support pharmicists in receiving and inputting prescriptions by phone.

– Correspond with patients regarding prescriptions and shipments.

• Wegmans Food Markets — Rochester, NY

Pharmacy Technician: October 2007 – July 2009 – Supervisor: Mark Murray

– Interact with customers to process and fill prescriptions.

– Work with team members to provide excellent and friendly customer service.

– Communicate with doctors and insurance companies to resolve conflicts.

• Wegmans Food Markets — Rochester, NY

Cake Artisan: October 2005 – October 2007 – Supervisor: Tom Kirkey

– Design and decorate special order cakes and desserts as requested by customers.

– Maintain a clean, healthy, and organized food preparation workspace.

Personal References

• Diane Thomas – 770 Bounty Drive #7006, Foster City, CA 94404

– (208) 720 - 2093

• Meg Daniels – 8451 Heritage Drive, Cincinnati, OH 45249

– (513) 314 - 8741

• Rebecca Schoonover – 101 Green Moor Way, Henrietta, NY 14467

– (607) 542 - 1867